OpenStudy (anonymous):

f(4) if f(x)=x^2-3x+3

OpenStudy (anonymous):

this kind of question basically ask you to calculate the value of the equation f with a value given of x. In this case, it is f(4)=4^2-3*4+3=16-12+3=7. Just put 4 in x's place and you are good to go :)

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

what about \[g(x)=2x-1 if g(3)

OpenStudy (anonymous):

you do the same thing, put 3 to where x is in this case, g(3)=2*3-1=6-1=5